An ebony log with volume 17.2 ft3 is submerged in water. What is the buoyant force on it (in lb)? (Enter the magnitude.)
1
Expert's answer
2019-04-15T10:12:27-0400

The buoyant force acting on the ebony log is equal to the weight of the displaced water:


"F_B = \rho_{water}V_{water}g."

Since the volume of displaced water by the ebony log, "V_{water}", is equal to the volume of the ebony log, "V_{log}", we can write:


"F_B = \rho_{water}V_{log}g,"

here, "F_B" is the buoyant force, "\rho_{water}" is the density of the water, "V_{log}" is the volume of the ebony log, "g" is the acceleration due to gravity.

Then, we get:


"F_B = 1000 \dfrac{kg}{m^3} \cdot 17.2 ft^3 \cdot \dfrac{0.02832 m^3}{1 ft^3} \cdot 9.8 \dfrac{m}{s^2} = 4774 N \cdot \dfrac{0.22481 lbf}{1 N} = 1073 lbf."

Answer:

"F_B = 1073 lbf."